OpenStudy (anonymous):

How many positive integers N are there such that the least common multiple of N and 1000 is 1000?

Parth (parthkohli):

This question reminds me of Euler's Totient Function.

Parth (parthkohli):

http://en.wikipedia.org/wiki/Euler's_totient_function There you go. You are only looking for \(\phi (1000)\).

terenzreignz (terenzreignz):

<Stares wide-eyed and with disbelief> O.O

terenzreignz (terenzreignz):

Too bad gourav isn't online right now :/ The prime factorisation for 1000 is \[\Large 2^3\cdot 5^3\] Now, if an integer were to have any prime factors OTHER THAN 2 or 5, then its Least Common multiple with 1000 would also have a prime factor other than 2 or 5, which means it can't be 1000. (Verify!) Now, we have established that any integer N must only have 2 or 5 for its prime factors. And of course, there can't be any powers of 2 or 5 greater than 3, or else, the LCM won't be 1000 (Verify!) So, we have \[\Large N = 2^p\cdot 5^q\qquad p,q\in\mathbb{Z}^+\qquad 0\le p,q\le 3\] This gives 16 possibilities. Namely, 1,2,4,8, 5,10,20,40, 25,50,100,200, 125,250,500,1000